dglwegf:

Arrange the functions in descending order, starting with the function that eventually has the greatest value and ending with the function that eventually has the least value. 4x + 9 9^x + 1 x^2 + 3 4x 9^x x^2

xxDeppressionxx:

I'll just assign a value of x and solve it so that I'll be able to arrange, them in order from greatest to least., I am assuming that x = 2., 4x + 9 4(2) + 9 = 8 + 9 = 17, 9^x + 1 9² + 1= 81 + 1 = 82, x² + 3 2² + 3 = 4 + 3 = 7, 4x 4(2) = 8, 9^x 9² = 81, x² 2² = 4, based on my assumption, the order from g